# Online Continual Graph Learning

### v1

@@ Abstract
The aim of Continual Learning (CL) is to learn new tasks incrementally while avoiding catastrophic forgetting. Online Continual Learning (OCL) specifically focuses on learning efficiently from a continuous stream of data with shifting distribution. While recent studies explore Continual Learning on graphs exploiting Graph Neural Networks (GNNs), only few of them focus on a streaming setting. Yet, many real-world graphs evolve over time, often requiring timely and online predictions. Current approaches, however, are not well aligned with the standard OCL setting, partly due to the lack of a clear definition of online Continual Learning on graphs. In this work, we propose a general formulation for online Continual Learning on graphs, emphasizing the efficiency requirements on batch processing over the graph topology, and providing a well-defined setting for systematic model evaluation. Finally, we introduce a set of benchmarks and report the performance of several methods in the CL literature, adapted to our setting.
Index Terms:
continual learning, online learning, graph neural networkI Introduction
In standard machine learning setups, models are at first trained in an environment where all training data is simultaneously available to the learning algorithm, and only at a later time model predictions are produced on new input data. Importantly, data observations are assumed to be mutually independent and identically distributed. Real-world environments however often generate data in chunks or streams which often undergo shifts in the data distribution or even variations in tasks to be solved. In turn, previously trained models may require expensive retraining or model reconfiguration to stay up to date. In this setting, Continual Learning (CL) [1, 2], lifelong learning [3] and incremental learning [4] are similar machine learning paradigms sharing the same goal of adapting models to incrementally learn as soon as new data and new tasks are presented.
Even further, in the online learning setting, training data points are collected sequentially, processed in real-time by the learning method, and immediately discarded [5, 6]. Such strict environments are found in monitoring and control problems [7, 8], such as traffic management, activity recognition, fraud detection, with other applications such as on data generated by optical sensors [9] or prediction of power production considering environmental conditions [10]. Online CL (OCL) is therefore an extremely challenging setting that requires models to quickly adapt to new conditions, allowing for anytime inference with minimal latency and without forgetting already acquired knowledge.
Research on CL extends to graph-structured data as well, a field often referred to as Continual Graph Learning (CGL) [11]. Indeed, many machine learning tasks involve graph representations of data, such as social networks, citation networks, biological systems and transaction networks. However, most graphs in the real world are not static: they continuously evolve over time experiencing the addition/removal of nodes and changes to their topology. Examples include social network growths, new publications in citation networks, and changes in road conditions [12, 13]. However, the literature has yet to formalize a rigorous online setting for CGL, and lacks a thorough analysis of its unique challenges. Among all, node-level predictions in graph machine learning often rely on past neighboring information, as seen in Graph Neural Networks (GNNs), leading to computational challenges as graphs densify over time.
In this paper, we provide four main contributions to address this gap.
@@
We formalize the Online Continual Graph Learning (OCGL) framework, establishing a foundation for CL on graphs in a node-level streaming environment. By doing so, we bridge the gap between OCL and CGL settings.
@@
We highlight the critical issue of neighborhood expansion, which could undermine the online computational efficiency due to the complexity of multi-hop aggregation in GNNs as the graph expands. To address it, we propose and discuss a first, simple and viable solution to ensure that models can efficiently scale as the graph evolves.
@@
We present a benchmarking environment for OCGL, encompassing six datasets, and evaluate several existing methods; as most of these methods are designed for offline CL, we suitably adapt them to operate under the OCGL setting.
@@
We conduct further empirical analyses, including ablation studies considering different settings and model configurations and evaluation over heterophilic graphs. Our findings reveal higher performance of replay-based methods that are tailored to preserve topological information in the considered setting.
We believe this work lays a strong foundation for future research to develop sound and more effective approaches for online continual graph learning.
@@ II Background and related works
Continual Learning. Depending on the type of shift in the data distribution, CL has been categorized into three main scenarios [14]: in task-incremental learning, the model sequentially learns distinct tasks, which requires availability of task identifiers to make predictions; class-incremental learning consists in classifying instances with an increasing number of classes, without task identifiers; finally, domain-incremental learning requires solving the same problem in different contexts. In the past, the main applications of CL included reinforcement learning [15, 16] and especially computer vision [17, 18, 19, 20, 21, 6], although most of the methods that have been developed to address these problem domains are not limited to these fields and can be used for a wide range of other machine learning tasks. Three main broad categories of CL approaches to mitigate forgetting have been proposed in the literature [2]: regularization methods, replay methods and architectural methods. Regularization methods [15, 22, 19, 20, 4] introduce additional loss terms to preserve important parameters to retain previously acquired knowledge. Replay methods [17, 18, 23, 5] use a memory buffer to store some representative samples from old tasks, to use them jointly with new samples to update model parameters. Architectural methods [24, 25, 26] avoid changes to model parameters either by gating mechanisms or by introducing new parameters, allowing the model to grow.
@@
Online Continual Learning. In the usual CL scenarios described above, data arrive one task at a time, allowing for offline training with multiple passes and shuffles over the data for the current task [2]. A more realistic and challenging scenario is that of Online Continual Learning (OCL) [5, 6, 27], where data arrive in small batches of only few samples, without the possibility for the model to store all the data for the current task, either for privacy reasons or memory limitations. In this setting the learning algorithm must be able to use efficiently the mini-batches that arrive in the non-stationary stream. Additionally, whereas for CL we assume to know the task boundaries, OCL can be performed in a boundary-agnostic setting, or task-free, potentially allowing for more varied distribution shifts compared to the three settings described before [28]. However, many CL methods are not suited to this setting and require modifications. An additional characteristic of OCL is the ability to perform anytime inference: the model should always be up-to-date and ready to make predictions online after each training batch, reacting quickly to distribution shifts [28].
@@
Learning on graphs. Graph Neural Networks (GNN) [29, 30, 31, 32] have emerged as the state-of-the-art approach for dealing with network data, generalizing convolution to graph structures. The core mechanism of most GNNs is message passing [33]: at each layer, the hidden embedding of each node is updated using information from its neighborhood as . Here AGGREGATE and UPDATE are differentiable functions specified by the particular model. Specifically, as at each step each node updates its embedding using the information (message) coming from its neighbors, after layers it will depend on its -hop neighborhood. Graph-based processing of temporal data has a relatively short history, primarily encompassing the study of temporal graphs [34, 35] and time series data [36, 37] with dedicated adaptation strategies to deal with evolving graphs [38] and benchmarks [39].
@@
Continual Graph Learning. In the last few years, researchers have started to develop CL strategies tailored to graph data [40], with applications such as recommender systems [41] and traffic prediction [42]. The main approaches resemble general CL strategies, with a particular focus on preserving topological information with a loss term on neighborhood aggregation parameters [12], or specific node selection policies to retain informative nodes as memory buffer [13]. Recently, a number of surveys have been published on the topic [43, 11, 44, 45], and a benchmark has been proposed [46]. Continual Graph Learning (CGL) possesses some peculiarities that differentiate it from other problem domains with independent samples, as graph structure requires careful consideration. Specifically, we can distinguish between graph-level CGL and node-level CGL [46]. In the first, the model needs to make predictions about entire graphs, and thus each sample in the training data is an independent graph. Thus, standard CL settings and methods can be applied with relative ease [47]. Instead, in node-level GCL predictions are made about the single nodes in usually one single graph. In this node-level CGL setting each new task therefore consists a new subgraph of the overall graph, with new classes or type of nodes. Specifically, the task subgraph arrives all at once, and offline training is performed on it. Here, an additional specification needs to be made about the availability of inter-task edges [45]: since message passing GNNs aggregate information from neighboring nodes, the presence of inter-task edges towards nodes of previous tasks would inevitably lead to the use of past data. In practice, inter-task edges are often kept, but without access to the labels from past tasks [44]. Finally, the addition of new nodes creates an additional source of backward interference, known as structural shift [48]: when evaluating the model on nodes belonging to past tasks, their neighborhood composition and topology will be changed compared to when the task was observed.
@@ III Online Continual Graph Learning
The OCL setting described previously has been explored in domains such as computer vision [6, 27] and sequences [49], but to the best of our knowledge, it has not yet been thoroughly studied for graph-structured data. Some papers on CGL consider a setting referred to as streaming [40, 50], yet the approaches can be categorized as offline CL as the streams consist of graph snapshots (often corresponding to entire tasks or nonetheless not classifiable as mini-batches), on which models are trained with multiple passes. While some CGL methods could in principle be used in an online fashion [43], the practical implications and constraints of the online setting for graph data have not been investigated yet. We highlight this gap in Table I and fill it with this paper.
We introduce Online Continual Graph Learning (OCGL), a new framework that ports CGL to the online problem setting. Specifically, OCGL is applicable to dynamic real-world scenarios such as social networks or recommender systems, where sudden distribution changes occur, and quick model adjustments are essential for making anytime predictions. We focus on the study of node classification, although the framework can be adapted to regression problems or edge-level tasks.
III-A Problem formulation
A growing network. We model the data associated with an OCGL problem as an evolving graph induced by a stream of nodes . At each time step , the monitored system is represented as a graph defined by node set , edge set , and a set of node attributes . Edge attributes, e.g., accounting for the strength of the connection or defining the type of node-node relations, can be considered likewise, however, they are not discussed here to ease the presentation. The graph nodes can be associated with class labels or target feature vectors collected in set to be predicted and/or used as training samples to learn the model. For simplicity, in this paper we focus on classification tasks, as most commonly done in CL. Graph is a snapshot of the temporal graph and, together with , collects all information available up to time step . At each time step a new node is acquired from the monitored system and populates the current snapshot graph . Specifically, tuple containing a new node index , associated node features , and a set of neighbors is presented and connected to graph according to the relations contained in . Finally, the target class label of node may or may not be acquired contextually to ; for instance, a prediction for node might be requested at time while the true class label is observed only at a later time.
The definition of the evolving graph is general, as it does not make assumptions on the distribution shifts happening in the node stream. This general, task-free node stream can be easily adapted or made more specific: while a real-world stream could be induced by a time-stamp on the nodes, this setting can be derived from any static graph by establishing an ordering on the nodes. The three CL scenarios of task-, class- and domain-incremental can thus be easily adapted to an online setting by ordering nodes by task, similarly to what is done in other domains [6, 27].
@@ III-A Problem formulation
Problem statement. The goal of OCGL is to incrementally update a prediction model , using only information associated with the current node and its neighbors of a given order within . Parameters have a predefined (maximum) dimension and model is trained incrementally as soon as new nodes and the associated labels are provided. New nodes are acquired either individually or in small mini-batches, slightly weakening the online setting as commonly done in the literature [5]. Importantly, in line with CL principles, the model must retain previously acquired knowledge: at time , the model may be asked to make predictions about , by using information from its neighbors in .
The key difference with standard CGL is that in OCGL, as in OCL, the model needs to adapt quickly in order to perform anytime predictions on the node stream [28]. This objective introduces specific efficiency constraints not present in the CGL literature; we refer the reader to Table I. The first requirement is to have small mini-batches of nodes: while in CGL settings training is done on a large task-specific subgraph, in OCGL this would be in contrast with the objective of anytime predictions. Due to their small size, mini-batches do not represent meaningful graph structures on their own. Consequently, OCGL relies on inter-batch edges. Another important requirement is that mini-batches are seen only once and, after training is performed, the mini-batch is discarded; a possible exception is the use of a fixed-size memory buffer that can store a limited number of samples. Finally, we require the training on each mini-batch to have bounded compute and memory budgets, independently from graph size at the current time step. This efficiency constraint poses non-negligible issues associated with reiterated message passing within multi-layer GNNs. We delve deeper into these complexities below, and we illustrate the difference of graph evolution compared to GCL in Figure 1.
@@
Neighborhood expansion problem. At each of the layers of the GNN, the network aggregates the embeddings of the neighboring nodes, thus requiring access to -hop neighborhoods of nodes. The size of the -hop neighborhood of a node however scales as where is the average degree. Moreover, is time-dependent and can increase as the graph grows: real-world graphs exhibit densification over time, with the number of edges following a power law of the number of nodes, and a shrinking of graph diameter over time [51, 52]. Depending on how well-connected a graph is, very few hops may be required to go from any node to almost any other one. Thus, large or would require processing a number of nodes in the order of the entire (growing) graph for each mini-batch, going against efficiency in our definition of OCGL. It is therefore of paramount importance to keep a relatively small or to introduce a strategy to deal with a large .
@@
To provide evidence of the neighborhood expansion phenomenon on four of the datasets considered in our experiments (see Section V for further details), in Figure 2 we plot the size of the -hop neighborhood of each mini-batch in the node stream inside the evolving graph, for some values of . The Reddit graph in particular is very well connected, with two hops containing the majority of the graph, and three hops covering practically all nodes. CoraFull and Arxiv show a more moderate neighborhood expansion, but still prohibitive. Lastly, two hops of the Amazon Computer graph covers, even in the worst case, about half of the nodes. Additional plots are shown in Appendix B, with the addition of the number of edges connecting the various hops, which can be used as a proxy of computational complexity.

### v2

@@ Abstract
Continual Learning (CL) aims to incrementally acquire new knowledge while mitigating catastrophic forgetting. Within this setting, Online Continual Learning (OCL) focuses on updating models promptly and incrementally from single or small batches of observations from a data stream. Extending OCL to graph-structured data is crucial, as many real-world networks evolve over time and require timely, online predictions. However, existing continual or streaming graph learning methods typically assume access to entire graph snapshots or multiple passes over tasks, violating the efficiency constraints of the online setting. To address this gap, we introduce the Online Continual Graph Learning (OCGL) setting, which formalizes node-level continual learning on evolving graphs under strict memory and computational budgets. OCGL defines how a model incrementally processes a stream of node-level information while maintaining anytime inference and respecting resource constraints. We further establish a comprehensive benchmark comprising seven datasets and nine CL strategies, suitably adapted to the OCGL setting, enabling a standardized evaluation setup. Finally, we present a minimalistic yet competitive baseline for OCGL, inspired by our benchmarking results, that achieves strong empirical performance with high efficiency.
I Introduction
In standard machine learning, models are trained once on a fixed dataset, assuming independent and identically distributed samples. Real-world environments, however, often generate data in chunks or streams, undergoing shifts in the data distribution or even variations in tasks to be solved, which requests often impractical and expensive periodic retraining. Continual Learning (CL) [1, 2] addresses this challenge by enabling models to learn incrementally while retaining past knowledge and without requiring all data to remain available.
In the more restrictive online learning setting, training data points are collected sequentially and must be processed by the learning method as soon as they appear and in real-time. Once processed, each sample is typically discarded leaving the learner with no access to past data [3, 4]. Such strict environments are found in monitoring and control problems [5, 6] where decisions must be made continuously and under limited resources. Building on this, Online Continual Learning (OCL) represents a particularly challenging setting in which models must rapidly adapt to evolving data streams using small batches that are observed only once. Additionally, they are expected to support anytime inference, meaning they are expected to generate accurate predictions at arbitrary points in time, even before training can be considered completed. At the same time, OCL systems must operate with minimal computational and memory cost, and mitigate forgetting of previously acquired knowledge.
Recently, CL has been extended to graph-structured data, giving rise to the field of Continual Graph Learning (CGL) [7]. Indeed, many machine learning tasks involve graph representations of data, such as social networks, citation networks, biological systems, and transaction networks. Additionally, most graphs in the real world are not static: they continuously evolve, experiencing the addition/removal of nodes and changes to their topology. Examples include the growth of social networks, the appearance of new publications in citation networks, and changes in road conditions [8, 9]. However, most existing CGL methods operate in offline and task-wise fashions, training on subgraph snapshots with multiple passes until convergence. Such settings fail to meet the core requirements of online learning – single-pass updates, limited budget, and anytime predictions – and overlook specific challenges of dynamic graphs. For instance, multi-hop message passing in Graph Neural Networks (GNNs) leads to unbounded computational growth as graphs densify due to the inclusion of new nodes and edges. We refer to this issue as the neighborhood expansion problem, which we analyze in this work. These additional constraints of the OCGL setting render most existing methods unsuitable for realistic, inherently evolving environments.
To address these shortcomings, we introduce the Online Continual Graph Learning (OCGL) setting, which unifies continual and online learning principles for node-level graph streams. We formalize the incremental acquisition of graph information under these constraints and analyze the performance of existing methods when adapted to this setting, along with their practical implications. Our main contributions are summarized as follows.
@@ I Introduction
We formalize OCGL as a principled setting bridging OCL and CGL, establishing a foundation for CL in environments characterized by data streams of node-level information.
@@
We highlight and discuss the neighborhood expansion problem, which OCGL introduces, and that can break the computation and memory requirements if not properly handled. We present a simple viable solution to address this problem.
@@
We design a benchmarking environment for OCGL, encompassing seven datasets and several existing methods from the literature, suitably adapted to operate within the OCGL setting. Our findings reveal higher performance of replay-based methods that are tailored to preserve topological information.
@@
We introduce LINEAR, a simple and lightweight method inspired by the observations drawn from the benchmark results. Despite its simplicity, LINEAR achieves competitive performance, and it constitutes a strong baseline to assess the quality of new OCGL methods.
We believe that this work lays a strong foundation for systematic progress in CL, fostering the development of sound and more effective approaches for OCGL, while LINEAR offers a strong and reliable reference for future research, setting a clear benchmark that new methods should meet or exceed to demonstrate meaningful progress.
@@ II-A Continual Learning
Depending on the type of shift in the data distribution, CL has been categorized into three main scenarios [10]: in task-incremental learning, the model sequentially learns distinct tasks, which requires availability of task identifiers to make predictions; class-incremental learning consists in classifying instances with an increasing number of classes, without task identifiers; finally, domain-incremental learning requires solving the same problem in different contexts. In the past, CL was mainly applied to reinforcement learning [11, 12] and computer vision [13, 14, 15, 16, 17, 4], but most of the methods that have been developed to address these problem domains can be used for a wide range of other machine learning tasks. CL approaches to mitigate forgetting fall into three general categories [2]: regularization, replay and architectural methods. Regularization methods [11, 18, 15, 16, 19] introduce additional loss terms to preserve important parameters to retain previously acquired knowledge. Replay methods [13, 14, 20, 3] use a memory buffer to store some representative samples from old tasks, to use them jointly with new samples to update model parameters. Architectural methods [21, 22, 23] avoid changes to model weights either by gating mechanisms or by introducing new parameters, allowing the model to grow.
II-B Online Continual Learning
@@
In the usual CL scenarios described above, data arrive one task at a time, allowing offline training with multiple passes and shuffles over the data for the current task [2]. Online Continual Learning (OCL) [3, 4, 24] addresses the more realistic case where data arrive in small batches of only few samples, without the possibility for the model to store all the data for the current task, either for privacy reasons or memory limitations. In this setting, the algorithm must efficiently learn from each mini-batch in a nonstationary stream. Additionally, whereas for CL we assume to know the task boundaries, OCL can be performed in a boundary-agnostic setting, or task-free, allowing for diverse distribution shifts [25]. However, many CL methods are not suited to this setting and require modifications. An additional characteristic of OCL is anytime inference: the model should always be up-to-date and ready to make predictions online after each training batch, reacting quickly to distribution shifts [25].
II-C Learning on graphs
@@
Graph Neural Networks (GNN) [26, 27, 28, 29] have emerged as the state-of-the-art approach for dealing with network data, generalizing convolution to graph structures. The core mechanism of most GNNs is message passing [30]: at each layer, the hidden embedding of each node is updated using information from its neighborhood as . Here AGGREGATE and UPDATE are differentiable functions specified by the particular model. Specifically, as at each step each node updates its embedding using the information (message) coming from its neighbors, after layers it will depend on its -hop neighborhood. Graph-based processing of temporal data has a relatively short history, primarily encompassing the study of temporal graphs [31, 32, 33, 34] and time series data [35, 36] with dedicated adaptation strategies to deal with evolving graphs [37] and benchmarks [38].
II-D Continual Graph Learning
@@
In the last few years, researchers have started to develop CL strategies tailored to graph data [39], with applications such as recommender systems [40] and traffic prediction [41]. Most Continual Graph Learning (CGL) methods adapt general CL strategies, focusing on preserving topological information with a loss term on neighborhood aggregation parameters [8], or specific node selection policies to replay informative nodes [9]. Recently, a number of surveys have been published on the topic [42, 7, 43, 44], and a benchmark has been proposed [45]. Importantly, CGL differs from other problem domains because of dependencies introduced by graph structure, requiring careful consideration. Specifically, we can distinguish between graph-level CGL and node-level CGL [45]. In graph-level CGL, each sample is an independent graph and standard CL methods apply directly [46], while node-level CGL performs predictions within a single evolving graph. In node-level CGL, each task consists of a new subgraph, for example with new classes of nodes. Specifically, the task subgraph arrives all at once, and offline training is performed on it. A key issue is the treatment of inter-task edges [44]: since GNNs aggregate neighbor information, edges to nodes from previous tasks may implicitly expose past data. In practice, inter-task edges are often kept, but without access to the labels from past tasks [43]. Finally, adding new nodes changes previous neighborhoods, causing structural shift [47], a source of backward interference.
@@ III Online Continual Graph Learning
The OCL setting has been explored in domains such as computer vision [4, 24] and sequences [48], but it has not been thoroughly investigated for graph-structured data. Some papers on CGL consider a setting referred to as streaming [39, 49], yet the approaches can be categorized as offline CL as the streams consist of graph snapshots, on which models are trained with multiple passes. While some CGL methods could be used in an online fashion [42], the practical implications and constraints of the online setting for graph data have not been investigated yet. Table I summarizes these differences and motivates the new setting introduced here.
We introduce Online Continual Graph Learning (OCGL), a new setting that ports CGL to the online problem setting. Specifically, OCGL is applicable to dynamic real-world scenarios such as social networks or recommender systems, where sudden distribution changes occur, and quick model adjustments are essential for anytime predictions. We focus on the study of node classification, although the setting can be adapted to regression problems or edge-level tasks. In this section we describe the general setting and its associated principles, requirements and challenges, while in Section V we will specify the instantiation of OCGL, such as the particular stream construction, used in our experiments.
III-A A growing network
We model the data associated with an OCGL problem as an evolving graph induced by a stream of nodes added in succession. At each time step , the graph snapshot is defined by node set , edge set , and a set of node attributes . Edge attributes can be considered as well, however they are not discussed here to ease the presentation. The graph nodes can be associated with class labels to be predicted and/or used as training samples to learn the model. At each time step a new node is added and connected to graph . Specifically, a tuple containing a new node index , associated node features , and a set of neighbors is presented and used to connect to graph according to the relations contained in . Finally, the target class label of node may or may not be acquired contextually to : for instance, a prediction for node might be requested at time while the true class label is observed only at a later time.
The formulation of the evolving graph is general, as it does not make assumptions on the distribution shifts happening in the node stream. It can be easily adapted or made more specific: while a real-world stream could be induced by a time-stamp on the nodes, this setting can be derived from any static graph by establishing an ordering on the nodes. The three CL scenarios of task-, class- and domain-incremental can thus be easily adapted to this online setting by ordering nodes by task, similarly to what is done in other domains [4, 24].
@@ III-B Problem statement
Given a model , the objective of OCGL is to incrementally update using only the information from the current node and its -hop neighborhood in , or using small mini-batches, slightly weakening the online setting as commonly done in the literature [3]. In line with CL principles, while adapting to the evolving stream the model must also retain previously acquired knowledge. The key difference with standard CGL is that in OCGL, as in OCL, the model needs to adapt quickly in order to perform anytime predictions on the node stream [25]. This objective introduces specific constraints (see also Table I):
New nodes arrive individually or in small mini-batches that are processed once; after training, each batch is discarded, possibly except for a limited replay buffer.
Differently from CGL, small mini-batches may not form meaningful self-contained subgraphs, hence inter-batch edges are naturally present and must be handled consistently.
Each update must have bounded computational and memory cost, regardless of the total size of .
The model must provide anytime inference, quickly adapting to new knowledge while retaining previously acquired one: after each update it can be queried on past nodes using the current graph state .
Satisfying these constraints guarantees efficiency and scalability as the graph grows to arbitrary size, but it poses non-negligible issues associated with reiterated message passing within multi-layer GNNs. We delve deeper into these complexities below, and we illustrate the difference of graph evolution compared to GCL in Figure 1.
III-C Neighborhood expansion problem
@@
At each of the layers of the GNN, the network aggregates the embeddings of neighboring nodes, thus requiring access to their -hop neighborhoods. The neighborhood size however scales as where is the average degree, which tends to increase as the graph grows: real-world graphs exhibit densification over time, with the number of edges following a power law of the number of nodes, and a shrinking of graph diameter over time [50, 51]. In many real graphs, only a few hops suffice to connect most nodes, so large or implies processing almost the entire graph per mini-batch, violating OCGL’s efficiency constraints. It is therefore of utmost importance to limit or to introduce a strategy to deal with . We note that neighborhood expansion is also present in CGL, but it is not in contrast with the requirements of that setting, and is thus not discussed in the literature.
@@
To illustrate this phenomenon, Figure 2 shows the size of the -hop neighborhood across mini-batches in the node stream for four datasets considered in our experiments (see Section V for further details). The Reddit graph in particular is very well connected, with two hops containing the majority of the graph. CoraFull and Arxiv show a more moderate neighborhood expansion, while two hops on the Amazon Computer graph cover about half of the nodes.
III-D Mini-batch acquisition